# When and How: Learning Identifiable Latent States for Nonstationary Time Series Forecasting

### v1

@@ E.1 Identification of Latent Domain Variables ${\mathbf{u}}_{t}$
where $\hat{p}({\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T})$ and $p({\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T})$ denote the estimated and ground-truth joint distributions, respectively; and $p({\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T})$ has transition matrix $\mathbf{A}$ and emission distribution $(\mu_{1},\cdots,\mu_{E})$, similarly for $\hat{p}({\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T})$. We consider any three consecutive observations with the order of ${\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3}$. According to A4, for any consecutive observations with the order of ${\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3}$, there are three different cases of the corresponding latent domain variables 1) $e_{1}=e_{2}=e_{3}$ 2) $e_{1}=e_{2}\neq e_{3}$ 3) $e_{1}\neq e_{2}=e_{3}$, and we discuss these cases, respectively. Note that the case of $e_{1}\neq e_{2}\neq e_{3}$ is impossible because there are at least 3 consecutive observations for each $e$.
Case 1. We consider the first case where any three consecutive or non-consecutive observations with the order of ${\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3}$ and the corresponding domain latent variables satisfy $e_{1}=e_{2}=e_{3}$.
$p({\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3})$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p({\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3},{\mathbf{z}}_{1}^{s},{\mathbf{z}}_{2}^{s},{\mathbf{z}}_{3}^{s},e_{1},e_{2},e_{3})d_{z_{1}^{s}}d_{z_{2}^{s}}d_{z_{3}^{s}}$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p({\mathbf{x}}_{1},e_{1},{\mathbf{z}}_{1}^{s})p({\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s}|{\mathbf{x}}_{1},e_{1},{\mathbf{z}}_{1}^{s})p({\mathbf{x}}_{3},e_{3},{\mathbf{z}}_{3}^{s}|{\mathbf{x}}_{1},e_{1},{\mathbf{z}}_{1}^{s},{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p({\mathbf{x}}_{1},e_{1},{\mathbf{z}}^{s}_{1}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}^{s}_{2})p({\mathbf{x}}_{2},e_{2},{\mathbf{z}}^{s}_{2})p({\mathbf{x}}_{3},e_{3},{\mathbf{z}}_{3}^{s}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p(e_{1}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s})p({\mathbf{z}}^{s}_{1}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s},e_{1})p({\mathbf{x}}_{1}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s},e_{1},{\mathbf{z}}^{s}_{1})p({\mathbf{x}}_{2},e_{2},{\mathbf{z}}^{s}_{2})p({\mathbf{x}}_{3},e_{3},{\mathbf{z}}_{3}^{s}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p(e_{1}|e_{2})p({\mathbf{z}}^{s}_{1}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{1}|e_{1},{\mathbf{z}}^{s}_{1})p({\mathbf{x}}_{2},e_{2},{\mathbf{z}}^{s}_{2})p({\mathbf{x}}_{3},e_{3},{\mathbf{z}}_{3}^{s}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$ (I)
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p(e_{1}|e_{2})p({\mathbf{z}}^{s}_{1}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{1}|e_{2},{\mathbf{z}}^{s}_{1})p({\mathbf{x}}_{2},e_{2},{\mathbf{z}}^{s}_{2})p({\mathbf{x}}_{3},e_{3},{\mathbf{z}}_{3}^{s}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$ (II)
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p(e_{1}|e_{2})p({\mathbf{z}}^{s}_{1}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{1}|e_{2},{\mathbf{z}}^{s}_{1})p({\mathbf{x}}_{2}|e_{2},{\mathbf{z}}^{s}_{2})p(e_{2})p({\mathbf{z}}^{s}_{2})p({\mathbf{x}}_{3},e_{3},{\mathbf{z}}_{3}^{s}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p(e_{1}|e_{2})p({\mathbf{z}}^{s}_{1}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{1}|e_{2},{\mathbf{z}}^{s}_{1})p({\mathbf{x}}_{2}|e_{2},{\mathbf{z}}^{s}_{2})p(e_{2})p({\mathbf{z}}^{s}_{2})p({\mathbf{z}}_{3}^{s}|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{3},e_{3},|{\mathbf{x}}_{2},e_{2},{\mathbf{z}}_{2}^{s},{\mathbf{z}}_{3}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p(e_{1}|e_{2})p({\mathbf{z}}^{s}_{1}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{1}|e_{2},{\mathbf{z}}^{s}_{1})p({\mathbf{x}}_{2}|e_{2},{\mathbf{z}}^{s}_{2})p(e_{2})p({\mathbf{z}}^{s}_{2})p({\mathbf{z}}_{3}^{s}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{3},e_{3}|e_{2},{\mathbf{z}}_{3}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p(e_{1}|e_{2})p({\mathbf{z}}^{s}_{1}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{1}|e_{2},{\mathbf{z}}^{s}_{1})p({\mathbf{x}}_{2}|e_{2},{\mathbf{z}}^{s}_{2})p(e_{2})p({\mathbf{z}}^{s}_{2})p({\mathbf{z}}_{3}^{s}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{3}|e_{2},{\mathbf{z}}_{3}^{s},e_{3})p(e_{3}|e_{2},{\mathbf{z}}_{3}^{s})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$
$=$ $\sum_{e_{1},e_{2},e_{3}}\int_{{\mathbf{z}}_{1}^{s}}\int_{{\mathbf{z}}_{2}^{s}}\int_{{\mathbf{z}}_{3}^{s}}p(e_{1}|e_{2})p({\mathbf{z}}^{s}_{1}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{1}|e_{2},{\mathbf{z}}^{s}_{1})p({\mathbf{x}}_{2}|e_{2},{\mathbf{z}}^{s}_{2})p(e_{2})p({\mathbf{z}}^{s}_{2})p({\mathbf{z}}_{3}^{s}|{\mathbf{z}}_{2}^{s})p({\mathbf{x}}_{3}|e_{2},{\mathbf{z}}_{3}^{s})p(e_{3}|e_{2})d_{{\mathbf{z}}_{1}^{s}}d_{{\mathbf{z}}_{2}^{s}}d_{{\mathbf{z}}_{3}^{s}}$
$=$ $\sum_{e_{2}}p(e_{2}){\left(\sum_{{\mathbf{e}}_{1}}\int_{{\mathbf{z}}^{s}_{1}}p(e_{1}|e_{2})p({\mathbf{z}}^{s}_{1}|{\mathbf{z}}^{s}_{2})p({\mathbf{x}}_{1}|e_{2},{\mathbf{z}}^{s}_{1})d_{{\mathbf{z}}_{1}^{s}}\right)}\left(\int_{{\mathbf{z}}_{2}^{s}}p({\mathbf{z}}^{s}_{2})p({\mathbf{x}}_{2}|e_{2},{\mathbf{z}}_{2}^{s})d_{{\mathbf{z}}_{2}^{s}}\right){\left(\sum_{e_{3}}\int_{{\mathbf{z}}^{s}_{3}}p({\mathbf{z}}_{3}^{s}|{\mathbf{z}}^{s}_{2})p(e_{3}|e_{2})p({\mathbf{x}}_{3}|e_{2},{\mathbf{z}}_{3}^{s})d_{{\mathbf{z}}_{3}^{s}}\right)}$
$=$ $\sum_{e_{2}}p(e_{2})\underbrace{\left(\sum_{{\mathbf{e}}_{1}}A_{e_{2},e_{1}}\cdot\mu_{e_{2}}\right)}_{\overline{\mu}_{e_{2}}}\mu_{e_{2}}\underbrace{\left(\sum_{e_{3}}A_{e_{2},e_{3}}\cdot\mu_{e_{2}}\right)}_{\dot{\mu}_{e_{2}}},$ (28)
@@
where we can obtain equation (II) from equation (I) by assuming that the value of $e_{1}$ and $e_{2}$ are the same. According A2 and A3, $\mathbf{A}$ is full rank and the probability measure $\mu_{1},\mu_{2},\cdots,\mu_{E}$ are linearly independent, the probability measure $\overline{\mu}_{e_{2}}=\sum_{{\mathbf{e}}_{1}}A_{e_{2},e_{1}}\cdot\mu_{e_{2}}$ are linearly independent and the probability measure $\dot{\mu}_{e_{2}}=\sum_{e_{3}}A_{e_{2},e_{3}}\cdot\mu_{e_{2}}$ are also linearly independent, Thus, applying Theorem 9 of (Allman et al., 2009), there exists a permutation $\sigma$ of $\{1,\cdots,E\}$, such that, $\forall i\in\{1,\cdots,E\}$:

### v4

@@ E.1 Identification of Latent Domain Variables ${\mathbf{u}}_{t}$
where $\hat{p}({\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T})$ and $p({\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T})$ denote the estimated and ground-truth joint distributions, respectively; and $p({\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T})$ has transition matrix $\mathbf{A}$ and emission distribution $(\mu_{1},\cdots,\mu_{E})$, similarly for $\hat{p}({\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T})$.
According to Theorem 1, since the nonstationary latent variables are block-wise identifiable, we can consider three consecutive nonstationary latent variables ${\mathbf{z}}_{1}^{e},{\mathbf{z}}_{2}^{e},{\mathbf{z}}_{3}^{e}$ and corresponding three discrete elements $e_{1},e_{2},e_{3}$.
$\begin{split}&p({\mathbf{z}}_{1}^{e},{\mathbf{z}}_{2}^{e},{\mathbf{z}}_{3}^{e})=\sum_{e_{1},e_{2},e_{3}}p({\mathbf{z}}_{1}^{e},{\mathbf{z}}_{2}^{e},{\mathbf{z}}_{3}^{e},e_{1},e_{2},e_{3})=\sum_{e_{1},e_{2},e_{3}}p(e_{2})p({\mathbf{z}}_{1}^{e},{\mathbf{z}}_{2}^{e},{\mathbf{z}}_{3}^{e},e_{1},e_{3}|e_{2})\\ =&\sum_{e_{1},e_{2},e_{3}}p(e_{2})p(z_{2}^{e}|e_{2})p({\mathbf{z}}_{1}^{e},{\mathbf{z}}_{3}^{e},e_{1},e_{3}|e_{2},{\mathbf{z}}_{2}^{e})=\sum_{e_{1},e_{2},e_{3}}p(e_{2})p(z_{2}^{e}|e_{2})p({\mathbf{z}}_{1}^{e},e_{1}|e_{2})p({\mathbf{z}}_{3}^{e},e_{3}|e_{2})\\ =&\sum_{e_{1},e_{2},e_{3}}p(e_{2})p(z_{2}^{e}|e_{2})p({\mathbf{z}}_{1}^{e}|e_{1})p(e_{1}|e_{2})p({\mathbf{z}}_{3}^{e}|e_{3})p(e_{3}|e_{2})\\ =&\sum_{e_{2}}p(e_{2})\underbrace{\Big{(}\sum_{e_{1}}p({\mathbf{z}}_{1}^{e}|e_{1})p(e_{1}|e_{2})\Big{)}}_{\bar{\mu}_{e_{2}}}\cdot\mu_{e_{2}}\cdot\underbrace{\Big{(}\sum_{e_{3}}p({\mathbf{z}}_{3}^{e}|e_{3})p(e_{3}|e_{2})\Big{)}}_{\dot{\mu}_{e_{2}}}.\end{split}$ (36)
@@
According A5 and A6, $\mathbf{A}$ is full rank and the probability measure $\mu_{1},\mu_{2},\cdots,\mu_{E}$ are linearly independent, the probability measure $\overline{\mu}_{e_{2}}=\sum_{{\mathbf{e}}_{1}}A_{e_{2},e_{1}}\cdot\mu_{e_{2}}$ are linearly independent and the probability measure $\dot{\mu}_{e_{2}}=\sum_{e_{3}}A_{e_{2},e_{3}}\cdot\mu_{e_{2}}$ are also linearly independent, Thus, applying Theorem 9 of (Allman et al.,, 2009), there exists a permutation $\sigma$ of $\{1,\cdots,E\}$, such that, $\forall i\in\{1,\cdots,E\}$: